OpenStudy (sarahc):

Find the fourth roots of 16(cos 200° + i sin 200°).

OpenStudy (michele_laino):

Please, you have to apply the De Moivre formula, I write your number as below: \[16*[\cos(200+k*360)+i*son(200+k*360)]\] then, your roots are: \[2[\cos(50+k*45)+i*\sin(50+k*45)]\] where k=0, 1, 2, 3 Please substitute those values for k and you will get all of your roots.

OpenStudy (michele_laino):

Sorry not k*45 but, k*90, namely th correct ormula is: \[2[\cos(50+k*90)+i*\sin(50+k*90)], k=0, 1,2 ,3\]

OpenStudy (michele_laino):

for example if we set k=0, we have the first root, namely: \[2[\cos50+i*\sin50]\] please do the same with k=1, 2, 3